OpenStudy (anonymous):

solve the given equation (1/x-1)+(2/x)=0

myininaya (myininaya):

\[\frac{1}{x-1}+\frac{2}{x}=\frac{x}{x} \cdot \frac{1}{x-1}+\frac{x-1}{x-1} \cdot \frac{2}{x}\] \[\frac{x+2(x-1)}{x(x-1)}=\frac{3x-2}{x(x-1)}\] now fractions are only zero where their top is zero (we don't want the bottom to be zero by the way) so set 3x-2=0 and that will be where your fraction is zero

OpenStudy (anonymous):

i dont get it